A box contains three two-sided coins of the same size and weight. Coin A is a standard

unbiased coin, while Coin B has two heads. Coin C is biased so that a tail is twice as likely to

occur as the head. A coin is randomly selected from the box, tossed and landed on "heads"

What is the probability that it is Coin C?


1
Expert's answer
2022-05-12T18:28:25-0400
"P(A)=P(B)=P(C)=\dfrac{1}{3}"

"P(H|A)=\dfrac{1}{2}, P(H|B)=1, P(H|C)=\dfrac{1}{3}"

"P(C|H)="

"=\dfrac{P(C)P(H|C)}{P(A)P(H|A)+P(B)P(H|B)+P(C)P(H|C)}"


"=\dfrac{\dfrac{1}{3}(\dfrac{1}{3})}{\dfrac{1}{3}(\dfrac{1}{2})+\dfrac{1}{3}(1)+\dfrac{1}{3}(\dfrac{1}{3})}=\dfrac{2}{11}"